1. Determine the molarity of a solution containing 10.6g of sodium carbonate dissolve in
a)250cm³ of distilled water
b)500cm³ of distilled water
1
Expert's answer
2020-04-24T13:33:42-0400

Solution:

Molarity or molar concentration is the number of moles of solute per liter of solution, which can be calculated using the following equation:

Molarity = mol solute / L of solution.


Molar mass of sodium carbonate = Mr(Na2CO3) = 105.9888 g/mol = 106 g/mol.

To find molaruty of Na2CO3 we need to find out how many moles of sodium carbonate are in solution. We can convert the mass of the solute to moles using the molar mass of sodium carbonate (Na2CO3), 106 g/mol:

Moles of Na2CO3 = Mass of Na2CO3 / Molar mass of Na2CO3

Moles of Na2CO3 = n(Na2CO3) = (10.6 g) / (106 g/mol) = 0.1 mol.


a) V(solution) = 250 cm³ = 0.25 L.

Molarity of Na2CO3 = n(Na2CO3) / V = (0.1 mol) / (0.25 L) = 0.4 M.

Molarity of Na2CO3 is 0.40 M.


Answer (a): Molarity of Na2CO3 is 0.40 M.


b) V(solution) = 500 cm³ = 0.50 L.

Molarity of Na2CO3 = n(Na2CO3) / V = (0.1 mol) / (0.50 L) = 0.20 M.

Molarity of Na2CO3 is 0.20 M.


Answer (b): Molarity of Na2CO3 is 0.20 M.